Help please with this review for physics

Help please with this review for physics
Answer #1

Solution 1) 20x106 x15 x105 20+15)x10 300x101 35x105 -8.57x10-F Ca = 8.57x10-6 +5x10-6 +10x10° 23.57x10-F -6 Hence the capaci

= 5x10-6x10 -50x10-6 C 10x106x10 100x10-C = 20x10-6x10 200 x 10 Q, = C.xv 15x10 x10 = 150x10-6C

50x10- 5x10-5 10V 02 100x10- 10x10-5 10V Q3 200x10 20x10-5 10V 150x10- 15x10-5 =10V

= 0.5 (23.57 x10-6)(10) 1178.5x10-*J Hence the energy stored in a capacitor is1178.5x10 J